# Meta-GGAs vs. Hybrid Functionals for Point Defects: The Best of Both Worlds Applied to Layered MnO2, NiO2 and KCoO2

### v2

@@ II.3 Layered $\mathrm{KCoO_{2}}$
Figure 9: r2SCAN+rVV10+U spin-resolved density of states per atom in (A) pristine KCoO2 with Co(III) ions with U=$\mathrm{U_{d}}$=1.8 eV, a single K-intercalated KCoO2 with (B) U=$\mathrm{U_{d}}$=0.0 eV, (C) U=$\mathrm{U_{d}}$=1.8 eV, (D) U=$\mathrm{U_{d}}$=0.9 eV. The more complicated pattern of in-gap states could reflect a Co(II) configuration different from a single unpaired electron in an eg state. Here we see the appearance of several localized occupied states on the defect Co(II) cation. The geometry was optimized in r2SCAN+rVV10+U+$\mathrm{U_{d}}$, which makes the interlayer spacing   5.98 Å. Like the d-states of Ni(IV), Co(III) has no electron in the eg state and completely filled t2g states. But, unlike the d-state of Ni(III), which has only one electron in the eg state and completely filled t2g states, Co(II) has four electrons in eg states, two in each channel, and partially filled t2g states, as shown in Fig. 11B. The DOS peaks appearing in up and down channels in the gap region of a K-intercalated KCoO2 are due to the eg electrons in the up and down channels of Co(II), respectively.
@@
We performed r2SCAN+rVV10+U+$\mathrm{U_{d}}$ calculations to study the defects in a K- intercalated KCoO2. First, we tried U=$\mathrm{U_{d}}$=0.0 eV to see how r2SCAN+rVV10 performs without U correction, and obtained a DOS as shown in Fig. 9B. The blue plot, which is the projected density of states (PDOS) of d-states of the Co(II) ion, shows a significant change in the d-state of Co(II) compared to the d-state of Co(III) ion. Our calculation shows the magnetic moment of the Co(II) ion around 3 , indicating three unpaired electrons in d-states. The presence of three unpaired electrons was previously seen in Refs. Sun et al. (2022); Chen et al. (2011); Kim et al. (2023); Bunting et al. (2018). We observed Co(II) d-state splitting, as shown in Fig. 11B. This explains the appearance of Co(II) d-states just above the VB maximum in both channels, absence of Co(II) d-states in the CB in the up channel and presence in the down channel. There is a breaking of spin symmetry in the DOS, due to the appearance of a spin moment on the defected Co ion. The phenomena like shifting of d-states to higher energy and spin symmetry breaking in CB are similar to those observed and explained before in Peng et al.’s work Peng et al. (2017) in the study of birnessite.
@@
As discussed in the Supplementary Material, we tried to determine the U value for the Co(II) ion by equating the HSE06+D3 magnetic moment to the r2SCAN+rVV10+U magnetic moment and obtained the U value for the Co(II) ion to be 0.91 eV. We used U=$\mathrm{U_{d}}$=0.91 eV and obtained a DOS as shown in Fig. 9D, which looks similar to Figs. 9B and 9C.
We tried to compare a full HSE06+D3 DOS with r2SCAN+rVV10+U+$\mathrm{U_{d}}$ for pristine KCoO2 and a K-intercalated KCoO2. We obtained the KCoO2 supercell by inserting a K layer for each CoO2 layer. Compared to the pristine CoO2 supercell, which contains 96 ions, the KCoO2 supercell has 32 extra K ions, and the K-intercalated KCoO2 supercell has 33 extra K ions. Therefore, one would expect a different lattice parameter for KCoO2 and a K-intercalated KCoO2 , which requires the full relaxation of the supercell. We attempted to optimize the structure using the HSE06+D3 method, but it is very challenging for these systems in terms of time and resource. After a month, we chose not to proceed further. This challange highlights one of the important reasons that motivated us to explore the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ method, highlighting the significance of the r2SCAN meta-GGA for defect studies.
@@
Fig. 11 shows our electronic configurations for Co(III) and Co(II) in octahedral or nearly octahedral coordination with oxygen. For Co(II), we have seven d electrons. There are three $\mathrm{t_{2g}}$ orbitals of each spin, and two $\mathrm{e_{g}}$ orbitals of each spin, and we can put no more than one electron into each spin orbital. Our output tells us the relative energy order of each occupied d state of each spin, that the net spin is nearly that of three spin-up electrons, and that the lowest-energy orbitals are spin-up while the highest-occupied orbital is spin-down, but not whether a given spin orbital is $\mathrm{t_{2g}}$ or $\mathrm{e_{g}}$. So we begin by putting three spin-up electrons into $\mathrm{t_{2g}}$ orbitals, to minimize the electronic repulsion. We assume that the $\mathrm{e_{g}}$ electrons prefer to be spin-up, to take advantage of an exchange interaction with the three up-spin $\mathrm{t_{2g}}$ orbitals, and that spin-up $\mathrm{e_{g}}$ levels will compete in energy with the spin-down $\mathrm{t_{2g}}$ levels. Then the configuration will be the one shown in Fig. 11 B and C, $\mathrm{(t_{2g}\uparrow)^{3}(t_{2g}\downarrow)^{2}(e_{g}\uparrow)^{1}(e_{g}\uparrow)^{1}}$. This result agrees with the $\mathrm{(t_{2g})^{5}(e_{g})^{2}}$ configuration found for Co(II) in $\mathrm{Li_{2}CoO_{2}}$ in Ref. Sun et al. (2022). This electronic configuration is also supported by the total charge density plots (see Fig. S5) of the four localized orbitals that appear in the band gap region of $\mathrm{K_{1.03}CoO_{2}}$.
@@
Fig. S4 B of the Supplementary Material shows the Jahn-Teller distorted tetrahedron around a defect Co(II) ion, with two of the Co(II)-O bonds about 10% longer than the remaining four. The angles are also distorted.
@@ III Conclusions
Here we demonstrate that the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ method can successfully describe defects in materials with proper values of U and $\mathrm{U_{d}}$. The method is much faster than hybrid functionals but equally accurate. Using the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ geometry enables the completion of hybrid functional calculations in a reasonable time and with fewer resources. This approach makes hybrid functional calculations feasible for larger systems with transition-metal ions.
The r2SCAN+rVV10+U+$\mathrm{U_{d}}$ method proposed here is not only faster but also as accurate as the already established hybrid functional method for defect studies. This method can expedite the study of defects in materials. The r2SCAN+rVV10+U+$\mathrm{U_{d}}$ method can also be employed in systems without defects where ions of the same species are in different OS.
@@
The r2SCAN+rVV10+U+$\mathrm{U_{d}}$ approach is successful, but it has limitations. It requires material-dependent parameters U, and those are typically available for only a few oxidation states of each transition metal. We have suggested a way to supplement the available values by choosing U to match the HSE06 magnetic moments, but that does not work when a moment is zero. A possibly more satisfactory approach would be to do r2SCAN+rVV10+U calculations for isolated transition-metal ions with integer M and half-integer M+1/2 electron numbers. Since the +U correction should be much less important for integer than for half-integer electron number, one can use U(M+1/2) for both M and M+1. Then U(M+1/2) could be found to satisfy the PPLB Perdew et al. (1982) condition E(M+1/2) = [E(M+1)+E(M)]/2, and U($\mathrm{M^{\prime}}$) could be found for all $\mathrm{M^{\prime}}$ by interpolation of U($\mathrm{M^{\prime}}$+1/2), yielding a set of oxidation-state-dependent values for each transition element. This possibility, which requires further thought and computation, is under consideration now in our group. Useful transfer of U from the free ion to the ion in a solid is not guaranteed, but the ratios of the U values for one element in different oxidation states might transfer better.
@@ IV Computational Details
First-principles calculations were performed with the projector-augmented wave method Blöchl (1994), implemented in the VASP code Kresse and Hafner (1994); Kresse and Joubert (1999). A 4 x 4 x 1 supercell was used to simulate defects in layered TMOs. For all supercell calculations, a 2 x 2 x 2 $\Gamma$-centered Monkhorst-Pack k mesh Monkhorst and Pack (1976) was used.
@@
For HSE (as for r2SCAN+rVV10+U+$\mathrm{U_{d}}$), corrections due to the periodic boundary condition and the supercell size have been added Peng et al. (2017); Kumagai and Oba (2014). For Fig. 2, we used an exact mixing parameter $\alpha$ of 0.22 in the hybrid functional HSE06 Peng et al. (2017); Krukau et al. (2006b), as determined from the Generalized Koopman’s Condition (GKC) method Lany and Zunger (2009); Lany (2011); Peng and Lany (2012); Peng et al. (2015). For all the remaining HSE06+D3 calculations, we used the standard exact exchange mixing parameter of 0.25. A plane-wave basis with an energy cutoff of 400 eV was employed.
@@ V Acknowledgments
This work was supported by the U.S. Department of Energy, Office of Science Basic Energy Sciences under Award No. DE-SC0023356.
VI Supplementary Material
Determining a U parameter from the HSE06 magnetic moment. Effect of the D3 dispersion correction on lattice constants and density of states. Jahn-Teller symmetry breaking of the CoO6 tetrahedron in a single K-intercalated KCoO2.
Supplementary Material
Determination of Hubbard U from Magnetic Moment
To determine a Hubbard U for some of our calculations, we equated the HSE06+D3 magnetic moment (MM) of ions inside the Wigner-Seitz sphere with the r2SCAN+rVV10+U’s MM. First, we relaxed the structure using the HSE06+D3 method and used this structure to calculate the HSE06+D3 MM. We then used the HSE06+D3 structure to do the r2SCAN+rVV10+U calculation, where we slowly varied the U value until r2SCAN+rVV10+U’s MM becomes equal to the HSE06+D3’s MM. We used smaller bilayer unit cells of transition metal oxides for U determination using this method. We used MnO2, KMnO2, KNiO2, and CaCoO2 to get U values for Mn(IV), Mn(III), Ni(III), and Co(II) ions, respectively. All calculations are performed using the VASP code, where we employed a Gamma-centered Monkhorst-pack grid of size 8 x 8 x 8 and a cut off energy of 400 eV for the plane-wave basis.
First, we used HSE06+D3 with an exact exchange mixing parameter ($\alpha$) of 0.22 to be consistent with Peng et. al.’s work Peng et al. (2017). We have summarized the calculated U value for these ions in Table S1 below.
HSE06($\alpha=0.22$)+D3 MM ($\mu_{B}$) r2SCAN+rVV10+U MM ($\mu_{B}$) U (eV)
MnO2 3.023 3.023 1.41
KMnO2 3.933 3.933 1.26
KNiO2 1.098 1.098 1.53
CaCoO2 2.663 2.663 0.55
Table S1: The U value determined by equating the magnetic moments (MM) of HSE06+D3 ($\alpha$ = 0.22) and r2SCAN+rVV10+U, using the HSE06+D3 ($\alpha$ = 0.22) geometry for both functionals.
We could not use this method to determine the U values for Ni(IV) and Co(III) ions using NiO2 and KCoO2 systems because the MM of Ni(IV) and Co(III) ions in NiO2 and KCoO2 is zero, largely independent of the U value, in the r2SCAN+rVV10+U calculation. So, we used the same U values for Ni(IV) and Co(III) as determined for Ni(III) and Co+4, respectively. We used the U values determined in Table S1, as well as U values from the literature Ding et al. (2021), and calculated lattice parameters for MnO2, NiO2, and CaCoO2 unit cells to see how r2SCAN+rVV10+U performs compared to HSE06+D3 with $\alpha$=0.22. The results are summarized in Table S2 below.
  System Method Lattice Parameters (Å) Lattice Angles (Degree)
a b c $\alpha$ $\beta$ $\gamma$
r2SCAN+rVV10 2.82 2.82 9.23 90 90 120
MnO2 r2SCAN+rVV10+U(1.41 eV) 2.83 2.83 9.22 90 90 120
HSE06($\alpha=0.22$)+D3 2.83 2.83 9.59 90 90 120
r2SCAN+rVV10 2.74 2.74 8.98 90 90 120
NiO2 r2SCAN+rVV10+U(1.53 eV) 2.73 2.73 9.00 90 90 120
HSE06($\alpha=0.22$)+D3 2.72 2.72 9.25 90 90 120
r2SCAN+rVV10 2.89 2.89 11.64 90 90 120
KCoO2 r2SCAN+rVV10+U(0.55 eV) 2.89 2.89 11.65 90 90 120
HSE06($\alpha=0.22$)+D3 2.86 2.86 11.21 90 90 120
Table S2: Lattice parameters and lattice angles comparison for r2SCAN+rVV10, r2SCAN+rVV10+U and HSE06($\alpha=0.22$)+D3 methods for the systems MnO2, NiO2 and KCoO2 .
We see that the lattice constants agree well for r2SCAN+rVV10 and r2SCAN+rVV10+U, but they disagree with HSE06+D3 ($\alpha$=0.22) lattice parameters, especially the c parameter. Probably this is because the D3 parameters were determined for HSE06 at $\alpha$ = 0.25, not 0.22.This motivated us to go back to full HSE06.
HSE06($\alpha=0.25$)+D3 MM ($\mu_{B}$) r2SCAN+rVV10+U MM ($\mu_{B}$) U (eV)
MnO2 3.036 3.036 1.71
KMnO2 3.947 3.947 1.58
KNiO2 1.163 1.163 2.06
CaCoO2 2.683 2.683 0.91
Table S3: The U value determined by equating the magnetic moments (MM) of HSE06+D3 ($\alpha$ = 0.25) and r2SCAN+rVV10+U, using the HSE06+D3 ($\alpha$ = 0.25) geometry for both functionals.
As above, we again determined U values by comparing the MM of r2SCAN+rVV10+U with the full HSE06+D3 ($\alpha$=0.25 ) functional. Table S3 below summarizes the result. The U values determined in the Table S3 are closer to the values in the literature, where optimal values for Mn, Ni, and Co ions for r2SCAN are 1.8 eV, 2.1 eV, and 1.8 eV, respectively Ding et al. (2021). The difference in oxidation states of ions could be the cause of the discrepancy. We used the above U values for the r2SCAN+rVV10+U functional and again compared the lattice parameters. The results are summarized in Table S4 below.
  System Method Lattice Parameters (Å) Lattice Angles (Degree)
a b c $\alpha$ $\beta$ $\gamma$
r2SCAN+rVV10 2.82 2.82 9.23 90 90 120
MnO2 r2SCAN+rVV10+U(1.71 eV) 2.84 2.84 9.11 90 90 120
HSE06($\alpha=0.25$)+D3 2.82 2.82 9.25 90 90 120
r2SCAN+rVV10 2.74 2.74 8.98 90 90 120
NiO2 r2SCAN+rVV10+U(2.06 eV) 2.73 2.73 8.98 90 90 120
HSE06($\alpha=0.25$)+D3 2.73 2.73 9.03 90 90 120
r2SCAN+rVV10 2.89 2.89 11.64 90 90 120
KCoO2 r2SCAN+rVV10+U(0.91 eV) 2.89 2.89 11.65 90 90 120
HSE06($\alpha=0.25$)+D3 2.87 2.87 11.61 90 90 120
Table S4: Lattice parameters and lattice angles comparison for r2SCAN+rVV10, r2SCAN+rVV10+U and HSE06($\alpha=0.25$)+D3 methods for the systems MnO2, NiO2 and KCoO2.
Figure S1: Spin-resolved density of states per atom using r2SCAN+rVV10+U+$\mathrm{U_{d}}$ (with U=$\mathrm{U_{d}}$=1.8 eV) geometry in a single K-intercalated MnO2 using (A) HSE06 (with $\alpha$=0.22) and (B) HSE06 (with $\alpha$=0.22)+D3.
Figure S2: (A) Mn(IV) - O bond lengths in pristine MnO2 and (B) defect Mn(III) - O bond lengths in a K-intercalated MnO2 ($\mathrm{K_{0.03}MnO_{2}}$) calculated using the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ (with U=$\mathrm{U_{d}}$=1.8 eV) functional, showing the Jahn-Teller symmetry breaking of the MnO6 octahedron. All lengths are in angstrom units. The red balls represent O ions and purple Mn ions. In the pristine materials, the three $\mathrm{t_{2g}\ (d_{xy},d_{yz},d_{zx})}$ space orbitals are equally populated, and the two $\mathrm{e_{g}\ (d_{x^{2}-y^{2}},d_{z^{2}})}$ spatial orbitals (whose lobes point from the metal cation to the 6 oxygens on the corners of the octahedron) are empty.
Figure S3: (A) Ni(IV) - O bond lengths in pristine NiO2 and (B) defect Ni(III) - O bond lengths in a K-intercalated MnO2 ($\mathrm{K_{0.03}NiO_{2}}$) calculated using the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ (with U=$\mathrm{U_{d}}$=2.1 eV) functional, showing the Jahn-Teller symmetry breaking of the NiO6 octahedron. All lengths are in angstrom units. The red balls represent O ions and silver Ni ions. In the pristine materials, the three $\mathrm{t_{2g}\ (d_{xy},d_{yz},d_{zx})}$ space orbitals are equally populated, and the two $\mathrm{e_{g}\ (d_{x^{2}-y^{2}},d_{z^{2}})}$ spatial orbitals (whose lobes point from the metal cation to the 6 oxygens on the corners of the octahedron) are empty.
Figure S4: (A) Co(III) - O bond lengths in pristine KCoO2 and (B) defect Co(II) - O bond lengths in a K-intercalated KCoO2 ($\mathrm{K_{1.03}CoO_{2}}$) calculated using the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ (with U=$\mathrm{U_{d}}$=1.8 eV) functional, showing the Jahn-Teller symmetry breaking of the CoO6 octahedron. All lengths are in angstrom units. The red balls represent O ions and blue Co ions. In the pristine materials, the three $\mathrm{t_{2g}\ (d_{xy},d_{yz},d_{zx})}$ space orbitals are equally populated, and the two $\mathrm{e_{g}\ (d_{x^{2}-y^{2}},d_{z^{2}})}$ spatial orbitals (whose lobes point from the metal cation to the 6 oxygens on the corners of the octahedron) are empty.
Figure S5: $\mathrm{r^{2}SCAN+rVV10+U+U_{d}\ (U=U_{d}=1.8\ eV)}$ based total charge density plots of four localized orbitals in the band gap region of $\mathrm{K_{1.03}CoO_{2}}$ around the defect Co(II) ion . The plots in the upper panel correspond to the two highest occupied orbitals in the spin up channel and clearly show them to be the $\mathrm{e_{g}}$ orbitals. The lower panel corresponds to the two highest occupied orbitals in the spin down channel and shows them to be the $\mathrm{t_{2g}}$ orbitals.

### v4

@@ II.3 Layered $\mathrm{KCoO_{2}}$
Figure 9: r2SCAN+rVV10+U spin-resolved density of states per atom in (A) pristine KCoO2 with Co(III) ions with U=$\mathrm{U_{d}}$=1.8 eV, a single K-intercalated KCoO2 with (B) U=$\mathrm{U_{d}}$=0.0 eV, (C) U=$\mathrm{U_{d}}$=1.8 eV, (D) U=$\mathrm{U_{d}}$=0.9 eV. The more complicated pattern of in-gap states could reflect a Co(II) configuration different from a single unpaired electron in an eg state. Here we see the appearance of several localized occupied states on the defect Co(II) cation. The geometry was optimized in r2SCAN+rVV10+U+$\mathrm{U_{d}}$, which makes the interlayer spacing $\approx$ 5.98 Å. In the fundamental gap of K-intercalated $\mathrm{KCoO_{2}}$, we see two localized $\mathrm{e_{g}}$uparrow electrons and two localized $\mathrm{t_{2g}}$downarrow electrons (Fig. 11).
@@
We performed r2SCAN+rVV10+U+$\mathrm{U_{d}}$ calculations to study the defects in a K- intercalated KCoO2. First, we tried U=$\mathrm{U_{d}}$=0.0 eV to see how r2SCAN+rVV10 performs without U correction, and obtained a DOS as shown in Fig. 9B. The blue plot, which is the projected density of states (PDOS) of d-states of the Co(II) ion, shows a significant change in the d-state of Co(II) compared to the d-state of Co(III) ion. Our calculation shows the magnetic moment of the Co(II) ion around 3 , indicating three unpaired electrons in d-states. The presence of three unpaired electrons was previously seen in Refs. Sun et al. (2022); Chen et al. (2011); Kim et al. (2023); Bunting et al. (2018). We observed Co(II) d-state splitting, as shown in Fig. 11B. This explains the appearance of Co(II) d-states just above the VB maximum in both channels, absence of Co(II) d-states in the CB in the up channel and presence in the down channel. There is a breaking of spin symmetry in the DOS, due to the appearance of a spin moment on the defected Co ion. The phenomena like shifting of d-states to higher energy and spin symmetry breaking in CB are similar to those observed and explained before in Peng et al.’s study Peng et al. (2017) of birnessite.
@@
As discussed in the SM SM- , we found the U value for Co(II) to be 0.91 eV. We used U=$\mathrm{U_{d}}$=0.91 eV and obtained a DOS as shown in Fig. 9D, which looks similar to Figs. 9B and 9C.
We tried to compare a full HSE06+D3 DOS with r2SCAN+rVV10+U+$\mathrm{U_{d}}$ for pristine KCoO2 and a K-intercalated KCoO2. We obtained the KCoO2 supercell by inserting a K layer for each CoO2 layer. Compared to the pristine CoO2 supercell, which contains 96 ions, the KCoO2 supercell has 32 extra K ions, and the K-intercalated KCoO2 supercell has 33 extra K ions. Therefore, one would expect a different lattice parameter for KCoO2 and a K-intercalated KCoO2 , which requires the full relaxation of the supercell. We attempted to optimize the structure using the HSE06+D3 method, but it is very challenging for these systems in terms of time and resource. After a month, we chose not to proceed further. This challenge highlights one of the important reasons that motivated us to explore the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ method, highlighting the significance of the r2SCAN meta-GGA for defect studies.
@@
Fig. 11 shows our electronic configurations for Co(III) and Co(II) in octahedral or nearly octahedral coordination with oxygen. For Co(II), we have seven d electrons. There are three $\mathrm{t_{2g}}$ orbitals of each spin, and two $\mathrm{e_{g}}$ orbitals of each spin, and we can put no more than one electron into each spin orbital. Our output tells us, from the relative energy order of each occupied d state of each spin, that the net spin is nearly that of three spin-up electrons, and that the lowest-energy orbitals are spin-up while the highest-occupied orbital is spin-down, but not whether a given spin orbital is $\mathrm{t_{2g}}$ or $\mathrm{e_{g}}$. So we begin by putting three spin-up electrons into $\mathrm{t_{2g}}$ orbitals, to minimize the electronic repulsion. We assume that the $\mathrm{e_{g}}$ electrons prefer to be spin-up, to take advantage of an exchange interaction with the three up-spin $\mathrm{t_{2g}}$ orbitals, and that spin-up $\mathrm{e_{g}}$ levels will compete in energy with the spin-down $\mathrm{t_{2g}}$ levels. Then the configuration will be the one shown in Fig. 11 B and C, $\mathrm{(t_{2g}\uparrow)^{3}(t_{2g}\downarrow)^{2}(e_{g}\uparrow)^{2}}$. This result agrees with the $\mathrm{(t_{2g})^{5}(e_{g})^{2}}$ configuration found for Co(II) in $\mathrm{Li_{2}CoO_{2}}$ in Ref. Sun et al. (2022). This electronic configuration is also supported by the total charge density plots (see Fig. S5) of the four localized orbitals that appear in the band gap region of $\mathrm{K_{1.03}CoO_{2}}$.
@@
Fig. S4 B of the SM SM- shows the Jahn-Teller distorted tetrahedron around a defect Co(II) ion, with two of the Co(II)-O bonds about 10% longer than the remaining four. The angles are also distorted.
@@ III Conclusions
Here we demonstrate that the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ method can successfully predict the existence of localized polaronic states near defects, with literature values of U and $\mathrm{U_{d}}$. The method is much faster than hybrid functionals but but still predicts localized polaronic states in band gaps. Using the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ geometry enables the completion of hybrid functional (or even G0W0) calculations in a reasonable time and with fewer resources. This approach makes hybrid functional calculations feasible for larger systems with transition-metal ions.
[…]
This method can expedite the study of defects in materials. The r2SCAN+rVV10+U+$\mathrm{U_{d}}$ method can also be employed in systems without defects where ions of the same species are in different OS.
@@
The r2SCAN+rVV10+U+$\mathrm{U_{d}}$ approach is successful, but it has limitations. It requires material-dependent parameters U, and those are typically available for only a few oxidation states of each transition metal. For cases where a literature value of U is unavailable, we have suggested (in the SM SM- ) a way to find a value by matching the HSE06 magnetic moment of the transition-metal ion (when it is non-zero). The HSE06 calculation could be made for a pristine material, avoiding large supercells. An interesting approach to U estimation, based on an ensemble of low-and high-spin states for a given transition metal ion and oxidation state, was suggested by Ref. Albavera-Mata et al. (2022).
@@ IV Computational Details
First-principles calculations were performed with the projector-augmented wave method Blöchl (1994), implemented in the VASP code Kresse and Hafner (1994); Kresse and Joubert (1999), version 5.4.4.. A 4 x 4 x 1 supercell was used to simulate defects in layered TMOs. For all supercell calculations, a 2 x 2 x 2 $\Gamma$-centered Monkhorst-Pack k mesh Monkhorst and Pack (1976) was used.
@@
For Fig. 2, we used an exact mixing parameter $\alpha$ of 0.22 in the hybrid functional HSE06 Peng et al. (2017); Krukau et al. (2006), as determined from the Generalized Koopman’s Condition (GKC) method Lany and Zunger (2009); Lany (2011); Peng and Lany (2012); Peng et al. (2015). For all the remaining HSE06+D3 calculations, we used the standard exact exchange mixing parameter of 0.25. A plane-wave basis with an energy cutoff of 400 eV was employed.
@@ V Acknowledgments
This work was supported by the U.S. Department of Energy, Office of Science Basic Energy Sciences under Award No. DE-SC0023356. We thank NERSC for computational resources.